# DiffHairCard: Auto Hair Card Extraction with Differentiable Rendering

### v1

@@ 4.5. Joint Optimization
After texture reduction, we perform another optimization with the visual losses and geometric regularization to match the collective visual appearance between the input strands =$\mathcal{H}=\{\mathcal{S}_{i}\}$ and our final cards $\{\mathcal{C}_{i}\}$ with uv-space strands $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ and tangent $\{\mathbf{t}_{i,j}^{\text{3D}}\}$.
@@
[…]
@@
[…]
@@
where $D(A,B)={2|A\cap B|}/{(|A|+|B|)}$ is the Sørensen-Dice coefficient to quantify the similarity between two masks, $A$ and $B$, where $|A|$ and $|B|$ are the total number of pixels in $A$ and $B$, and $|A\cap B|$ is the common area between $A$ and $B$. For rendering the binary mask image, we set all pixels to be one in a all-zero background. Note that we do not need the tangents $\mathbf{t}_{i,j}^{\text{3D}}$ for mask rendering.
@@
As our final step, after optimization, we rasterize the UV-space strands and tangents into RGB-format textures to generate the final tangent, depth, and alpha maps, as shown in Fig. 4. Additionally, we bake an ambient occlusion (AO) texture to enhance visual detail and store local lighting information on the hair cards. Note that the $z$ offset along normal direction in Eq. 2 is not optimized, which is essential for accurate AO baking.
@@ 4.6. Extensions
To enhance both the visual fidelity of our output and the practical usability of our pipeline, we introduce the following two additional features, which can be tuned on optionally.
@@
To simulate hair volume and improve visual fidelity from multiple viewing angles, we introduce crossed hair cards during generation. Specifically, for each cluster, we generate a pair of hair cards placed at a 90-degree angle to each other, creating the illusion of volumetric hair using flat geometry, as shown in Fig. 5. In practice, during the card generation stage (Sec. 4.2), we simply create an additional card perpendicular to the primary card at each segment by rotating both $\mathbf{n}_{k,j}$ and $\mathbf{t}_{k,j}$ over 90 degrees, giving a crossed configuration.
@@
Since hair cards alone may not fully cover the scalp under certain camera angles or lighting conditions, particularly near the roots, we construct a hair cap, a base layer of geometry and texture applied to the scalp that simulates the appearance of short, dense hair, so, it can efficiently represent root fuzz and base hair density with minimal computational cost. To create the hair cap mesh and its associated texture, we begin by projecting all hair roots onto their nearest faces on the head mesh $\mathcal{M}^{\text{head}}$. We then extract all faces containing at least one hair root, along with their one-ring neighboring faces to form the hair cap mesh. In order to prevent z-fighting during rendering, we slightly extrude the vertices of the cap mesh along their normal directions by a small distance $\epsilon^{\text{cap}}$. Finally, for each strand, we extract the first sub-segment with a fixed arc-length of $\epsilon^{\text{root}}$ and bake the sub-segment onto the scalp texture. Specifically, we bake the tangent, alpha, and AO into three texture channels as shown in Fig. 6.
@@ 5. Results
We evaluate our pipeline on a diverse set of hairstyles, including short, bangs, straight, ponytail, bun, wavy, curly, and coily, from MetaHuman (Epic Games, 2021) and dataset in previous high-fidelity hair reconstruction work (Shen et al., 2023). All hair models are down-sampled to use 40K hair strands, each with $n^{s}=32$ samples. Following the hair card texture settings used in MetaHuman, we set our output texture atlas size to $2048\times 2048$, which accommodates $n^{t}=4\times 8=32$ individual hair card textures, each at a resolution of $512\times 256$ by default. The card orientation optimization uses a screen size $128\times 128$, while joint optimization uses $256\times 256$ to capture strand-level details. To assess the visual similarity to the input strand-based model, we use three averaged metrics, PSNR $\uparrow$ for the shading evaluation, LPIPS $\downarrow$ for perceptual similarity, especially high-frequency hair details, and coverage error $\downarrow$ for the hair silhouette. The average values of these metrics are computed over 12 uniformly distributed viewpoints around the hair.
Ablation study on # card and # texture
Fig. 11 presents a comparison between our method and Unreal Engine’s automatic hair card generator (Epic Games, 2025) under varying number of card and texture budgets. We evaluate results using 200, 100, and 50 cards, targeting low-resolution LoD hair generation. For textures, we test with 16 and 32 textures, following the industrial best practices for texture budgeting. As expected, visual similarity to the reference strand-based model decreases as the card and texture budgets are reduced. Therefore, key to creating a good hair model lies in finding a balance between visual quality and geometric efficiency. Nevertheless, our method consistently achieves higher visual fidelity in all three metrics, even when using fewer textures and only 50 cards.
Ablation study on losses
We further conduct an ablation study on all loss and regularization terms on a straight hairstyle from MetaHuman (Epic Games, 2021) with 100 hair cards and 32 textures. As shown in Fig. 7, each component is necessary to achieve consistent and accurate results. For the rest of experiments, we adopt two sets of loss weights for different hair types. For straight hair, we use $\lambda^{\text{tangent}}=10$, $\lambda^{\text{depth}}=10$, $\lambda^{\text{dice}}=5$, $\lambda^{\text{match}}=3$, and $\lambda^{\text{collision}}=1\times 10^{5}$. For curly hair, we use a different configuration as $\lambda^{\text{tangent}}=5$, $\lambda^{\text{depth}}=15$, $\lambda^{\text{dice}}=3$, $\lambda^{\text{match}}=3$, and $\lambda^{\text{collision}}=1\times 10^{5}$, to encourage more curl strands.
@@
Comparison with optimizing texture directly
We validate the necessity of our explicit strand representation by comparing it to a conventional differentiable rendering pipeline that directly optimizes hair card textures using Nvdiffrast (Laine et al., 2020). In this experiment, we provide a fixed hair card geometry along with tangent images from eight view directions. We then use Nvdiffrast to optimize the texture mapped to the card so that its rendered appearance matches the given tangent images as closely as possible from those views. Although Nvdiffrast can produce hair textures with distinct shape features when optimizing for a single view direction, extending the optimization to multiple views leads to noisy and inconsistent results. Introducing a smoothness loss can eliminate noise at the cost of losing high-frequency details, as shown in Fig. 8. More importantly, this texture-based approach lacks the spatial structure for computing ambient occlusion accurately, as hair-to-hair spatial relationships are not preserved during texture optimization.
@@
Comparison with Unreal Engine’s auto-generated and manual-crafted cards
We compare our results with the output of Unreal Engine’s built-in automatic hair card generation tool (Epic Games, 2025) and artist-crafted hair cards from MetaHuman (Epic Games, 2021). We compare with the lowest LoD hair card model from MetaHuman, which only contains 351 cards and 20 textures. As shown in Fig. 9, using the same set of 351 hair cards and 20 textures, our method achieves better visual fidelity to the input strand-based model than both Unreal Engine’s auto-generated and manual-crafted cards. While our higher PSNR partly benefits from more accurate AO textures, as we preserve the displacement during optimization, the lower LPIPS and coverage errors clearly demonstrate that our results more faithfully preserve hair occupancy and shape.
@@
17.71/0.279/0.044
17.78/0.209/0.038
19.99/0.140/0.021
Figure 9. Comparison with Unreal Engine’s auto-generated and manual-crafted cards. Our method outperforms the other two in all metrics with the same 351 cards and 32 textures.
@@
Table 1. Statistics about card # used in both our method and Unreal Engine’s auto cards, as well as the corresponding visual metric statistics, PSNR $\uparrow$, LPIPS $\downarrow$, and coverage error $\downarrow$. Crossed card is indicated by ×$\times 2$.
Unreal Engine (Epic Games, 2025) Ours
Model Fig. # Card # PSNR $\uparrow$ LPIPS $\downarrow$ Coverage $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Coverage $\downarrow$
Ponytail 5 100$\times$2 20.91 0.123 0.022 22.70 0.095 0.010
Short 6 100$\times$2 22.34 0.112 0.014 23.51 0.100 0.010
Bun 9 351 17.78 0.209 0.038 19.99 0.140 0.021
Straight 11 100 17.39 0.230 0.038 18.21 0.176 0.034
Curly 12 400 18.54 0.153 0.038 20.89 0.138 0.016
Bangs 12 100 18.09 0.218 0.032 20.47 0.154 0.018
Blowout 12 200 17.16 0.200 0.038 20.44 0.162 0.017
Wavy 12 200 16.82 0.228 0.041 19.89 0.166 0.016
Coily 13 400$\times$2 17.90 0.174 0.032 19.09 0.131 0.022
Real-time demo
Fig. 10 demonstrates that our hair cards can be used directly in Unreal Engine with real-time rendering and simulation. Specifically, we generate one guide hair per card for simulation and drive the deformation of the hair cards using linear blend skinning. Simulating a hair card model with 512 cards takes only 0.12 ms per frame. Please refer to the supplemental video.
Figure 10. Simulating our output hair card model in Unreal Engine, taking only 0.12 ms per frame.

### v2

@@ 4.5. Joint Optimization
After texture reduction, we perform another optimization with the visual losses and geometric regularization to match the collective visual appearance between the input strands $\mathcal{H}=\{\mathcal{S}_{i}\}$ and our final cards $\{\mathcal{C}_{i}\}$ with uv-space strands $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ and tangent $\{\mathbf{t}_{i,j}^{\text{3D}}\}$.
@@
where quad strip position $\mathbf{p}_{k,j}^{\pm}$, strand texture coordinates $\mathbf{s}_{i,j}^{\text{uv}}$, tangent $\mathbf{t}^{\text{3D}}_{i,j}$, and per-strand width $w_{i}$ are all included as decision variables. To ensure high output quality, we combine visual loss terms $v\in\{\text{tangent},~\text{depth},~\text{dice}\}$ and geometry regularization terms $g\in$ $\{\text{match},~\text{collision}\}$, which are detailed below.
@@
(5) $L^{v}=\int_{\mathbb{S}^{2}}\mathsf{MSE}\bigl{(}\mathcal{I}^{v}(\mathcal{H},\omega),\mathcal{I}^{v}(\langle\mathcal{C}_{k},\mathbf{s}_{i,j}^{\text{uv}},\mathbf{t}_{i,j}^{\text{3D}},w_{i}\rangle,\omega)\bigr{)}d\omega,$
@@
where $D(A,B)={2|A\cap B|}/{(|A|+|B|)}$ is the Sørensen-Dice coefficient to quantify the similarity between two masks, $A$ and $B$, where $|A|$ and $|B|$ are the total number of pixels in $A$ and $B$, and $|A\cap B|$ is the common area between $A$ and $B$. For rendering the binary mask image, we set all pixels to one in a zero background. Note that we do not need the tangents $\mathbf{t}_{i,j}^{\text{3D}}$ for mask rendering.
@@
As our final step, after optimization, we rasterize the UV-space strands and tangents into RGB-format textures to generate the final tangent, depth, and alpha maps, as shown in Fig. 4. Additionally, we bake an ambient occlusion (AO) texture to enhance visual detail and store local lighting information on the hair cards. We convert explicit hair card texture into 3D tubes in UV-space and illuminate them with a directional light along the z-direction. Next, we perform standard offline ray-traced AO baking, computing surface color contributions after multiple light bounces. Finally, the resulting shaded surface is projected back onto the card plane to create the AO texture. Note that while it is possible to also optimize the $z$ offset along the normal direction in Eq. 3, we found in practice that keeping it fixed during optimization can provide sufficient results.
@@ 4.6. Extensions
To enhance both the visual fidelity of our output and the practical usability of our pipeline, we introduce two additional optional features.
@@
A common limitation of using billboards/cards is that when the view direction becomes nearly parallel to the card plane, the card appears invisible. To simulate hair volume and improve visual fidelity from multiple viewing angles, we introduce crossed hair cards during generation. Specifically, for each cluster, we generate a pair of hair cards placed at a 90-degree angle to each other, creating the illusion of volumetric hair using flat geometry, as shown in Fig. 5. In practice, during the card generation stage (Sec. 4.2), we simply create an additional card perpendicular to the primary card at each segment by rotating both $\mathbf{n}_{k,j}$ and $\mathbf{t}_{k,j}$ over 90 degrees, giving a crossed configuration.
@@
Since hair cards alone may not fully cover the scalp under certain camera angles or lighting conditions, particularly near the roots, we construct a hair cap, a base layer of geometry and texture applied to the scalp that simulates the appearance of short, dense hair, so it can efficiently represent root fuzz and base hair density with minimal computational cost. To create the hair cap mesh and its associated texture, we begin by projecting all hair roots onto their nearest faces on the head mesh $\mathcal{M}^{\text{head}}$. We then extract all faces containing at least one hair root, along with their one-ring neighboring faces, to form the hair cap mesh. In order to prevent z-fighting during rendering, we slightly extrude the vertices of the cap mesh along their normal directions by a small distance $\epsilon^{\text{cap}}$. Finally, for each strand, we extract the first sub-segment with a fixed arc-length of $\epsilon^{\text{root}}$ and bake the sub-segment onto the scalp texture. Specifically, we bake the tangent, alpha, and AO into three texture channels as shown in Fig. 6.
5. Implementation Details
This section provides a detailed description of the implementation of our rendering pipeline employed in both the optimization stage and the final results presentation.
Optimization Stage
After reconstructing all strand control points using Eq. 3, the strands are rasterized into camera-oriented quad strips with a user-defined strand width through Nvdiffrast (Laine et al., 2020). Since our optimization is guided by loss functions defined over tangent, depth, and coverage, no hair BCSDF is employed for shading. Instead, the renderer produces these G-buffer maps with the corresponding attributes for evaluating the loss function.
Rendering Final Results
To ensure high-quality rendering results, we employ an offline path tracing pipeline for final image generation. Hair appearance is modeled using the classical formulation of hair BCSDF (Marschner et al., 2003), and the local frames are reconstructed from tangent and depth textures during BCSDF sampling and evaluation. Upon ray–card intersection, the alpha texture is used to determine the strand occupancy at the texel level. Hair-hair occlusion is separated into two stages: inter-card occlusion is directly resolved via ray tracing, while intra-card strand occlusion is approximated using our precomputed AO texture.
Table 1. Statistics about card # used in both our method and Unreal Engine’s auto cards, as well as the corresponding visual metric statistics, PSNR $\uparrow$, LPIPS $\downarrow$, coverage error $\downarrow$, and chamfer distance (CD) $\downarrow$. Crossed card is indicated by $\times 2$.
Unreal Engine (Epic Games, 2025) Ours
Model Fig. # Card # PSNR $\uparrow$ LPIPS $\downarrow$ Coverage $\downarrow$ CD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Coverage $\downarrow$ CD $\downarrow$
[…]
Straight 7 100 17.39 0.230 0.038 0.005 18.21 0.176 0.034 0.005
Curly 9 400 18.54 0.153 0.038 0.005 20.89 0.138 0.016 0.005
Bangs 9 100 18.09 0.218 0.032 0.006 20.47 0.154 0.018 0.006
Blowout 9 200 17.16 0.200 0.038 0.005 20.44 0.162 0.017 0.005
Wavy 9 200 16.82 0.228 0.041 0.006 19.89 0.166 0.016 0.006
Bun 11 351 17.78 0.209 0.038 0.004 19.99 0.140 0.021 0.004
Braid 13 400 23.25 0.086 0.010 0.003 24.90 0.077 0.007 0.003
Fringe 14 100 18.47 0.166 0.036 0.007 19.24 0.158 0.029 0.008
[…]
@@ 6. Results
We evaluate our pipeline on a diverse set of hairstyles, including short, bangs, straight, ponytail, bun, wavy, curly, and coily, from MetaHuman (Epic Games, 2021) and a dataset in previous high-fidelity hair reconstruction work (Shen et al., 2023). All hair models are down-sampled to use 40K hair strands, each with $n^{s}=32$ samples. Following the hair card texture settings used in MetaHuman, we set our output texture atlas size to $2048\times 2048$, which accommodates $n^{t}=4\times 8=32$ individual hair card textures, each at a resolution of $512\times 256$ by default. The card orientation optimization uses a screen size $128\times 128$, while the joint optimization uses $256\times 256$ to capture strand-level details. For card geometry fitting, we sample 20 candidate normals evenly distributed along a positive half-circle. During optimization, we uniformly sample 64 viewpoints on the unit sphere, oriented toward the origin, to balance between coverage and computational efficiency. We set $n^{q}$ to 31 for standard cards and 15 for cross cards, except for Afro, we use 48. To assess the visual similarity to the input strand-based model, we use three averaged metrics, as presented in Table 1. PSNR $\uparrow$ for the shading evaluation, LPIPS $\downarrow$ for perceptual similarity, especially high-frequency hair details, and coverage error $\downarrow$ for the hair silhouette. We also use chamfer distance $\downarrow$ to evaluate the geometric accuracy. The average values of these metrics are computed over 12 uniformly distributed viewpoints around the hair.
Strands
Unreal Engine’s cards
Ours 16 textures
Ours 32 textures
Ours 64 textures
Reference
50/32, 16.97/0.270/0.048
50/16, 17.66/0.203/0.045
50/32, 17.99/0.184/0.40
N/A
100/64, 17.60/0.214/0.035
100/16, 18.03/0.192/0.035
100/32, 18.21/0.176/0.034
100/64, 18.51/0.165/0.029
200/64, 17.70/0.200/0.032
200/16, 18.99/0.159/0.027
200/32, 19.03/0.150/0.025
200/64, 19.38/0.138/0.021
[…]
Ablation Study on Card # and Texture #
Fig. 7 presents a comparison between our method and Unreal Engine’s automatic hair card generator (Epic Games, 2025) under varying numbers of cards and texture budgets. We evaluate results using 200, 100, and 50 cards, targeting low-resolution LoD hair generation. For textures, we test with 16 and 32 textures, following the industrial best practices for texture budgeting. As expected, visual similarity to the reference strand-based model decreases as the card and texture budgets are reduced. Therefore, the key to creating a good hair model lies in finding a balance between visual quality and geometric efficiency. Nevertheless, our method consistently achieves higher visual fidelity in all three metrics, even when using fewer textures and only 50 cards.
@@
Ablation Study on Losses
We further conduct an ablation study on all loss and regularization terms on a straight hairstyle from MetaHuman (Epic Games, 2021) with 100 hair cards and 32 textures. As shown in Fig. 8, each component is necessary to achieve consistent and accurate results. Unfortunately, due to the subpixel nature of hair strands and their complicated geometries, traditional image-based metrics such as MSE and PSNR cannot fully capture the perceptual quality of hair. For instance, omitting tangent loss often results in unnatural zigzag strands. Similarly, allowing cards to intersect with the head may seem minor in visual metrics, but it causes noticeable artifacts during simulation. For the rest of the experiments, we adopt two sets of loss weights for different hair types. For straight hair, we use $\lambda^{\text{tangent}}=10$, $\lambda^{\text{depth}}=10$, $\lambda^{\text{dice}}=5$, $\lambda^{\text{match}}=3$, and $\lambda^{\text{collision}}=1\times 10^{5}$. For curly hair, we use a different configuration as $\lambda^{\text{tangent}}=5$, $\lambda^{\text{depth}}=15$, $\lambda^{\text{dice}}=3$, $\lambda^{\text{match}}=3$, and $\lambda^{\text{collision}}=1\times 10^{5}$, to encourage more curl strands. Note that, due to the inherent geometric differences between straight and curly hair, we adopt two sets of configurations that work well across all examples. We believe this choice does not compromise the generalizability of our method.
Strands
UE cards
Our cards without cap
Our cards
with cap
Short
22.34/0.112/0.014
23.34/0.100/0.011
23.51/0.100/0.010
Bangs
18.09/0.218/0.032
20.38/0.154/0.019
20.47/0.154/0.018
Curly
18.54/0.153/0.038
20.75/0.138/0.018 20.89/0.138/0.016
Refer to caption Refer to caption Refer to caption Refer to caption
Blowout 17.16/0.200/0.038 20.40/0.162/0.017 20.44/0.162/0.017
Refer to caption Refer to caption Refer to caption Refer to caption
Wavy 16.82/0.228/0.041 19.89/0.166/0.018 19.89/0.166/0.016
Figure 9. Our results outperform UE cards over a variety of hair styles, including short, bangs, wavy, curly, and blowout, regardless of whether with a hair cap or not.
[…]
Ablation Study on Hair Caps
Our results outperform Unreal Engine’s cards over a variety of hair styles, including short, bangs, wavy, curly, and blowout, as shown in Fig. 9. The hair cap primarily improves scalp coverage and hairline, but does not affect overall hair silhouette. As for the geometry, hair cards inherently involve a trade-off between visual fidelity and geometric accuracy. This work primarily aims to reproduce the overall appearance within a constrained geometric budget. Hair geometry that is barely visible contributes little to appearance, so precise geometric accuracy in those regions is less critical. For this reason, our pipeline is appearance-driven, leveraging differentiable rendering to optimize hair cards for visual similarity rather than exact geometry reproduction.
Strands
Init w/ UE cards
Ours
17.27/0.232/0.046
19.30/0.158/0.027
Figure 10. Ablation study on card initialization. Using UE cards as initialization degrades the visual quality.
[…]
Strands
Manual cards
UE auto cards
Our cards
17.71/0.279/0.044
17.78/0.209/0.038
19.99/0.140/0.021
Figure 11. Comparison with Unreal Engine’s auto-generated and manually crafted cards. Our method outperforms the other two in all metrics with the same 351 cards and 32 textures.
[…]
@@
23.25/0.086/0.010
24.90/0.077/0.007
Figure 13. Our method applies equally well to complex structures such as knots and braids.
@@
Strands
UE cards
Our cards
Fringe
18.47/0.166/0.036
19.24/0.158/0.029
Figure 14. Hair card extraction for strand-based hairs from multi-view reconstruction (Zakharov et al., 2024).
[…]
Ablation Study on Card Initialization
We conduct an ablation study on card orientation optimization on Bun hairstyle from MetaHuman (Epic Games, 2021) with 200 hair cards and 32 textures. As shown in Fig. 10, using cards generated from the UE automatic hair card generator as the initial shape for our optimization results in a degradation of visual quality, while our card geometry initialization leads to better optimized results.
Comparison with UE Auto-generated and Manual-crafted Cards
We compare our results with the output of UE built-in automatic hair card generation tool (Epic Games, 2025) and artist-crafted hair cards from MetaHuman (Epic Games, 2021). We compare with the lowest LoD hair card model from MetaHuman, which only contains 351 cards and 20 textures. As shown in Fig. 11, using the same set of 351 hair cards and 20 textures, our method achieves better visual fidelity to the input strand-based model than both Unreal Engine auto-generated and manual-crafted cards. While our higher PSNR partly benefits from more accurate AO textures, as we preserve the displacement during optimization, the lower LPIPS and coverage errors clearly demonstrate that our results more faithfully preserve hair occupancy and shape.
Strands
Our cards
Figure 15. Comparison between simulating and rendering the input strand-based model and our output hair card model in Unreal Engine (Epic Games, 2021).
[…]
Comparison with Optimizing Texture Directly
We validate the necessity of our explicit strand representation by comparing it with a conventional differentiable rendering pipeline that directly optimizes hair card textures using Nvdiffrast (Laine et al., 2020). In this experiment, we provide a fixed hair card geometry along with tangent images from eight view directions. We then use Nvdiffrast to optimize the texture mapped to the card so that its rendered appearance matches the given tangent images as closely as possible from those views. Although Nvdiffrast can produce hair textures with distinct shape features when optimizing for a single view direction, extending the optimization to multiple views leads to noisy and inconsistent results. Introducing a smoothness loss can eliminate noise at the cost of losing high-frequency details, as shown in Fig. 12. More importantly, this texture-based approach lacks the spatial structure for computing ambient occlusion accurately, as hair-to-hair spatial relationships are not preserved during texture optimization.
Strands
Ours, 20 cards
Ours, 50 cards
Ours, 100 cards
Ours, 200 cards
20/16, 17.65/0.245/0.041
50/16, 17.66/0.203/0.045
100/16, 18.03/0.192/0.035
200/16, 18.99/0.159/0.027
[…]
Braid Styles
Fig. 2 demonstrates that buns can be grouped based on their geometric and spatial similarity, a strategy that applies equally well to complex structures such as knots and braids. Fig. 13 shows another example of braided hairstyles, where our method can produce hair cards that reproduce the braided style with a small number of hair cards.
Strand-based Hairs from Multi-view Reconstruction
To further validate the scope of application of our method, we test our method on the strand model generated using the multi-view reconstruction (Zakharov et al., 2024). As shown in  Fig. 14, our method can automatically extract hair cards from a reconstructed strand-based fringe hairstyle.
Comparison with Strand Model in Game Engine
Fig. 15 demonstrates that our hair cards can be used directly in Unreal Engine with real-time rendering and simulation. Specifically, we generate one guide hair per card for simulation and drive the deformation of the hair cards using linear blend skinning (LBS) as mentioned in (Hsu et al., 2024). The strand-based model with approximately 50,000 strands requires 2.4 ms per frame for rendering, whereas our card model only takes 0.8 ms for 512 hair cards. Please refer to the supplemental video.
@@
Strands
Unreal Engine’s 800 cards
[…]
Ours, $00×\times2$ cards
800/32, 17.90/0.174/0.032
[…]
$00×\times2$/32, 21.45/0.093/0.011
[…]